OpenStudy (anonymous):

how to implicitly differentiate v=(4/3)(pi)(D/2)^3 for a related rates problem in calc

OpenStudy (anonymous):

if you are differentiating with respect to time ti would be... hold on is this \[\frac{4\pi}{3}(\frac{D}{2})^3\]?

myininaya (myininaya):

\[v'=\frac{4}{3} \cdot \pi \cdot (\frac{1}{2})^3\cdot 3 \cdot D^{2} D'\]

OpenStudy (anonymous):

a bit of arithmetic to clean up there for sure.

OpenStudy (anonymous):

ya, thanks guys. i got it.